NOTE: Staple your assignment together before turning it in. I. Suppose the industry demand function for bacon takes the following form where Qd is the quantity of bacon demanded, P, is the price of bacon, P, is the price of sausage, and Pe is the price of eggs If Pb-10, P, and Pe4 what is the quantity of bacon demanded? b. Graph the demand curve for bacon when P, 1 and Pe4. What is the slope of this demand curve? c. Suppose the industry supply function for bacon is as follows: where Qb5 is the quantity of bacon supplied, Pb is the price of bacon, and Ph is the price of hogs. Suppose also that Ph- 4. Graph the supply curve for bacon d. Solve for the equilibrium quantity and price of bacon at the prevailing prices for eggs, sausages, and hogs. Show this equilibrium in a graph. At this equilibrium price and quantity, solve for the own price elasticity of demand for bacon? Is the demand curve at this point elastic, inelastic, or unit elastic? f. At this equilibrium price and quantity, solve for the cross-price elasticity of demand for bacon with respect to the price of sausage. Are bacon and sausage complements or substitutes? g. At this equilibrium price and quantity, solve for the own price elasticity of supply. Is supply at this point elastic, inelastic, or unit elastic?I was able to answer 1a and 1b -- but I'm stuck on C, D and E (you can disregard F and G unless you wanna answer them)
Answer #1

(c) When Ph = 4,

Qs = Pb - (0.5 x 4) = Pb - 2

Data table used:

Pb Qs = Pb - 2
0 -2
1 -1
2 0
3 1
4 2
5 3
6 4
7 5
8 6
9 7
10 8
11 9
12 10
13 11
14 12
15 13
16 14
17 15
18 16
19 17
20 18

Graph:

20 18 16 14 Supply 12- 10 8 4 18 8 -2 Quantity13

(d) Plugging in given values in demand function,

Qd = 48 - 4Pb

In equilibrium, Qd = Qs.

48 - 4Pb = Pb - 2

5Pb = 50

Pb = 10

Q = 10 - 2 = 8

In following graph, equilibrium is at point A with price P0 (= 10) and quantity Q0 (= 8).

12 -supply 10 A PO Dem and 8 4 2 Q0 -5 17 28 39 50

(e) From demand function,

Own price elasticity = (dQd/dPb) x (Pb/Qd) = -4 x (10/8) = -5

Since absolute value of own price elasticity is higher than 1, demand is elastic.